OpenStudy (anonymous):

Give the exact value of each of the following: csc 5π/6

OpenStudy (anonymous):

ok you have the \[csc\frac{5pi}{6}\] or \[5*\pi=5*180=900\] \[5\pi/6=900/6=150\] so you need to find \[\csc(150)\]

OpenStudy (anonymous):

\[csc(150)=csc(90+60)\] \[csc(90+60)=sec(60)\] \[sec(60)=2\]

OpenStudy (anonymous):

I do the same thing with this problem: sec 5π/6

OpenStudy (anonymous):

Lol why do you do that? Let a = 5Pi/6 so I don't have to type as much csc a = (sin a)^-1 csc a = (sin Pi/6)^-1 since sine is the same for the angle Pi/6 and 5Pi/6. = (1/2)^-1 = 2 For sec a, sec a = (cos a)^-1 sec a = -(cos Pi/6)^-1 since cosine is the negative for the two angles sec a = (1/sqrt3)^-1 sec a = - sqrt 3